ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 192, 2012

Ardizzoni, A. and A. Pavarin Preantipodes for dual quasibialgebras

281

Balandraud, É. Erratum to: “An addition theorem and maximal zero-sum free sets in ℤ/pℤ”

1009

Barros, A., J. Silva and P. Sousa Rotational linear Weingarten surfaces into the Euclidean sphere

819

Bell, J. P. and A. Smoktunowicz Rings of differential operators on curves

297

Bergweiler, W. On the packing dimension of the Julia set and the escaping set of an entire function

449

Biró, A. A duality relation for certain triple products of automorphic forms

587

Breuillard, E., B. Green, R. Guralnick and T. Tao Strongly dense free subgroups of semisimple algebraic groups

347

Casolo, C., S. Dolfi and E. Jabara Finite groups whose noncentral class sizes have the same p-part for some prime p

197

Castro, A. L. and R. Montgomery Spatial curve singularities and the Monster/Semple tower

381

Cheng, X., Z. Shen and Y. Tian A class of Einstein (α, β)-metrics  221

Ciperiani, M. and D. Krashen Relative Brauer groups of genus 1 curves

921

Climenhaga, V. and D. J. Thompson Intrinsic ergodicity beyond specification: β-shifts, S-gap shifts, and their factors

785

Daws, M., R. Haydon, T. Schlumprecht and S. White Shift invariant preduals of ℓ_{1}(ℤ)

541

Dietmann, R. and C. Elsholtz Hilbert cubes in progression-free sets and in the set of squares

59

Dolfi, S., G. Malle and G. Navarro The finite groups with no real p-elements

831

Dolfi, S. See C. Casolo

Dupont, C. Large entropy measures for endomorphisms of ℂℙ^{k}  505

Elsholtz, C. See R. Dietmann

Fridman, B. L. and D. Ma Testing holomorphy on curves  535

Friedman, S. D. and R. Honzik A definable failure of the singular cardinal hypothesis 719

Galindo, C. Clifford theory for graded fusion categories  841

Ginzburg, D., D. Jiang and D. Soudry On correspondences between certain automorphic forms on Sp_{4n}(𝔸) and S̃p_{2n}(𝔸)

951

Green, B. See E. Breuillard Guralnick, R. See E. Breuillard Haydon, R. See M. Daws Honzik, R. See S. D. Friedman

Hoyt, C. Good gradings of basic Lie superalgebras  251
Ilinca, L. and J. Kahn The number of 3-SAT functions  869

Jabara, E. See C. Casolo Jiang, D. See D. Ginzburg Juhász, I. and M. Magidor On the maximal resolvability of mono-

tonically normal spaces

637

Kahn, J. See L. Ilinca Karasev, R. N. and F. V. Petrov Partitions of nonzero elements

of a finite field into pairs

143

Krashen, D. See M. Ciperiani

Letzter, E. S. On prime ideals of Noetherian skew power series rings  67

Lima, V. and N. L. Randrianarivony Property (β) and uniform quotient maps

311

Lötscher, R. Essential dimension of involutions and subalgebras  325

Ma, D. See B. L. Fridman Magidor, M. See I. Juhász Malle, G. See S. Dolfi

Marberg, E. Generalized involution models for wreath products  157
Meier, M. An infinitary probability logic for type spaces  1

Montgomery, R. See A. L. Castro Naor, A. and T. Tao Scale-oblivious metric fragmentation and the

nonlinear Dvoretzky theorem

489

Navarro, G. See S. Dolfi Osȩkowski, A. Sharp weak-type inequalities for Hilbert transform

and Riesz projection

429

Pavarin, A. See A. Ardizzoni Petrov, F. V. See R. N. Karasev Pfeifle, J., V. Pilaud and F. Santos Polytopality and Cartesian

products of graphs

121

Pilaud, V. See J. Pfeifle Randrianarivony, N. L. See V. Lima Reyes, M. L. Obstructing extensions of the functor Spec to noncom-

mutative rings

667

Roychowdhury, M. K. Quantization dimension and temperature function for bi-Lipschitz mappings

473

Santos, F. See J. Pfeifle Schlumprecht, T. See M. Daws Shen, Z. See X. Cheng

Shin, S. W. Automorphic Plancherel density theorem  83

Silva, J. See A. Barros Smoktunowicz, A. See J. P. Bell Soudry, D. See D. Ginzburg Sousa, P. See A. Barros Tao, T. See A. Naor Tao, T. See E. Breuillard Thompson, D. J. See V. Climenhaga Tian, Y. See X. Cheng van Handel, R. On the exchange of intersection and supremum of

σ-fields in filtering theory

763

White, S. See M. Daws Ye, S. A unified approach to the plus-construction, Bousfield localiza-

tion, Moore spaces and zero-in-the-spectrum examples 699